OpenStudy (anonymous):

Please help, cos2x + cos3x sinx = 1/2

OpenStudy (anonymous):

cos^2x = 1-sin^2x. So, substitute and see what you get.

OpenStudy (mertsj):

\[\cos ^{2}x+\cos ^{3}xsinx=\frac{1}{2}\]

OpenStudy (mertsj):

Is that the equation?

OpenStudy (mertsj):

Or is it \[\cos2x+\cos3xsinx=\frac{1}{2}\]

OpenStudy (anonymous):

Your version may be the right one as I can't see how the "former one" I interpreted could hold true.